BHC Corporation just issued a $1,000 par value bond with a 7 percent yield to maturity,...

  1. BHC Corporation just issued a $1,000 par value bond with a 7 percent yield to maturity, twenty years to maturity, with an 8 percent semi-annual coupon rate. 6 points
  1. What is the price of the BHC Corporate bond?
  1. If market interest rates are constant, what will the price of the BHC Corporate bond be in three years?
  1. If market interest rates rise to 10 percent, what will the price of the BHC Corporate bond be in three years?
Answer #1

a]

Price of a bond is the present value of its cash flows. The cash flows are the coupon payments and the face value receivable on maturity

Price of bond is calculated using PV function in Excel :

rate = 7%/2 (Semiannual YTM of bonds = Annual YTM / 2)

nper = 20 * 2 (20 years to maturity with 2 semiannual coupon payments each year))

pmt = 1000 * 8% / 2 (semiannual coupon payment = face value * coupon rate / 2)

fv = 1000 (face value receivable on maturity)

PV is calculated to be $1,106.78

b]

Price of a bond is the present value of its cash flows. The cash flows are the coupon payments and the face value receivable on maturity

Price of bond is calculated using PV function in Excel :

rate = 7%/2 (Semiannual YTM of bonds = Annual YTM / 2)

nper = 17 * 2 (17 years to maturity with 2 semiannual coupon payments each year))

pmt = 1000 * 8% / 2 (semiannual coupon payment = face value * coupon rate / 2)

fv = 1000 (face value receivable on maturity)

PV is calculated to be $1,098.50

c]

Price of a bond is the present value of its cash flows. The cash flows are the coupon payments and the face value receivable on maturity

Price of bond is calculated using PV function in Excel :

rate = 10%/2 (Semiannual YTM of bonds = Annual YTM / 2)

nper = 17 * 2 (17 years to maturity with 2 semiannual coupon payments each year))

pmt = 1000 * 8% / 2 (semiannual coupon payment = face value * coupon rate / 2)

fv = 1000 (face value receivable on maturity)

PV is calculated to be $838.07